OpenStudy (anonymous):

Please help... Show all working out!! The volume of ice-cream in a cone is half the volume of the cone. The cone has a 3cm radius and 14cm height. What is the depth of the ice-cream correct to two decimal places?

OpenStudy (anonymous):

At first, tell me the formula of volume of a cone

OpenStudy (anonymous):

1/3 x Area of base x height

OpenStudy (anonymous):

Yep, and area of a base=pi r^2

OpenStudy (anonymous):

Now, find the volume of a cone

OpenStudy (anonymous):

Yeah, so it's 42pi .... Therefore, the volume of the ice-cream is 21pi.

OpenStudy (anonymous):

But how am I supposed to find the depth of the ice-cream?

OpenStudy (anonymous):

I guess thats wrong

OpenStudy (anonymous):

volume of a cone = pi (r)^2 h = pi(3)^2 * 14=126pi

OpenStudy (anonymous):

But the volume of a cone is 1/3 x area of base x height, so you need to divide 126pi by 3...

OpenStudy (anonymous):

So you get 42pi as the volume of the cone, and 21pi as the volume of the ice-cream.

OpenStudy (anonymous):

oh ya

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

However, I am unsure how to calculate the depth of the ice-cream

OpenStudy (anonymous):

no